OpenStudy (anonymous):

convert the following degree measure to radian measure 47 degrees type answer as a simplified fraction. type an exact answer in terms of pi

Parth (parthkohli):

For converting into radians, you divide by 180 and then put \(\pi\) in front of your result.

OpenStudy (anonymous):

so it would be pi 47/180

Parth (parthkohli):

Examples,\[360^{\circ} \implies {360 \over 180}\pi \implies 2\pi\]\[180 \implies {180 \over 180}\pi \implies \pi\]

Parth (parthkohli):

Actually, it'd be\[{47 \over 180}\pi \], yes!

OpenStudy (anonymous):

thank you.